# Normalized AOPC: Fixing Misleading Faithfulness Metrics for Feature Attribution Explainability

### v1

@@ Abstract
Deep neural network predictions are notoriously difficult to interpret. Feature attribution methods aim to explain these predictions by identifying the contribution of each input feature. Faithfulness, often evaluated using the area over the perturbation curve (AOPC), reflects feature attributions’ accuracy in describing the internal mechanisms of deep neural networks. However, many studies rely on AOPC to compare faithfulness across different models, which we show can lead to false conclusions about models’ faithfulness. Specifically, we find that AOPC is sensitive to variations in the model, resulting in unreliable cross-model comparisons. Moreover, AOPC scores are difficult to interpret in isolation without knowing the model-specific lower and upper limits. To address these issues, we propose a normalization approach, Normalized AOPC (NAOPC), enabling consistent cross-model evaluations and more meaningful interpretation of individual scores. Our experiments demonstrate that this normalization can radically change AOPC results, questioning the conclusions of earlier studies and offering a more robust framework for assessing feature attribution faithfulness.
Code — https://github.com/JoakimEdin/naopc
Introduction
@@
Deep neural networks are often described as “black boxes” due to the difficulty in understanding their reasoning processes (Wei et al. 2022). This lack of interpretability can hinder their adoption in critical applications where trust is paramount (Lipton 2018). For instance, if a diagnostic model predicts meningitis without an explanation, a physician confident in an influenza diagnosis might incorrectly dismiss the model’s prediction as an error.
@@
Feature attribution methods attempt to address this issue by quantifying each input feature’s contribution to a model’s output (Danilevsky et al. 2020). In the meningitis example, such a method might identify “fever” and “stiff neck” as important features, potentially convincing the physician to reconsider the diagnosis. For these methods to be reliable, they must faithfully represent the model’s underlying reasoning process, avoiding misleading interpretations (Jacovi and Goldberg 2020).
The Area Over the Perturbation Curve (AOPC) has become a standard metric for approximating faithfulness, with two main variants: sufficiency and comprehensiveness (DeYoung et al. 2020; Lyu, Apidianaki, and Callison-Burch 2024). Numerous studies have used AOPC scores to compare faithfulness across different models, often concluding that certain models (e.g., adversarially robust models) are more faithful than others (Bhalla, Srinivas, and Lakkaraju 2023; Li et al. 2023; Chrysostomou and Aletras 2022; Liu et al. 2022; Resck, Raimundo, and Poco 2024; Chen and Ji 2020; Sekhon et al. 2023; Chrysostomou and Aletras 2021; Nielsen et al. 2023).
However, our research uncovers a major flaw in this approach: AOPC scores are not directly comparable across different models and are difficult to interpret in isolation. We demonstrate that the minimum and maximum possible AOPC scores vary significantly across different models, even when they are evaluated on the same input. These model-specific lower and upper AOPC limits are caused by the models’ reasoning process, specifically how many features they rely on and their feature interactions. To illustrate this issue, we found that for one model, the upper limit AOPC score was 0.3, while for another model on the same dataset, it was 0.8. This discrepancy makes the AOPC scores produced by these models incomparable. Additionally, without knowing a model’s upper and lower limits, an AOPC score cannot be properly interpreted. For instance, is an AOPC score of 0.25 high? It is if the upper bound is 0.3, but not if it is 0.8.
@@
We demonstrate the model-specific nature of AOPC score limits, which renders cross-model comparisons and isolated score interpretations problematic.
@@
We show empirically how NAOPC alters faithfulness rankings, highlighting the need to re-evaluate previous conclusions about model faithfulness.
Problem formulation
Area Over the Perturbation Curve (AOPC) measures the change in model output as input features are sequentially perturbed (Samek et al. 2016). The perturbation can either remove, insert, or replace a feature with some pre-defined value. The final score is the average output change across all perturbation steps. Formally, the AOPC is calculated as follows:
where is a model, is an input vector with number of features, is the order to perturb the input features, and is the perturbation function that removes, inserts or replaces the features in that are in . AOPC is used to calculate sufficiency and comprehensiveness as follows.
Comprehensiveness estimates the faithfulness of feature attribution scores by perturbing the input features in decreasing order, starting from the feature with the highest score (DeYoung et al. 2020). A high comprehensiveness indicates that the features that are the highest ranked, according to the feature attributions, are important for the model’s output (i.e., higher is better). Comprehensiveness is calculated as follows:
@@
Sufficiency perturbs the input features in increasing order, starting from the feature with the lowest score (DeYoung et al. 2020). In other words, sufficiency is comprehensiveness when flipping the feature ordering . A low sufficiency indicates that the lowest ranked features, according to the feature attributions, are irrelevant to the model output (i.e., lower is better). Sufficiency is calculated as follows:
@@
Models influence AOPC scores
Ideally, sufficiency and comprehensiveness should only measure the feature attribution’s faithfulness. However, in this section, we demonstrate that a model’s reasoning process heavily influences these AOPC metrics. Specifically, we show that 1) the more features a model relies on, the worse the sufficiency and comprehensiveness scores, and 2) a model’s features interactions impact the best possible comprehensiveness and sufficiency scores.
@@
The two models use the same architecture, but their parameter values differ. relies heavily on feature , while relies on more features. Different training strategies, data, or randomness can cause such model differences (Hase, Xie, and Bansal 2021).
Given an input vector , both models output . Since the models are linear, we can calculate the ground truth feature attributions by multiplying each input feature by its parameter. For , this results in for , and for . As these attributions represent the ground truths, one might expect the models to yield equal sufficiency and comprehensiveness scores. However, this is not the case. As shown in table 1, achieves drastically worse sufficiency and comprehensiveness scores than simply because it relies on more features. While relying on fewer features could be a desirable model property, it should be measured using entropy instead of influencing the faithfulness evaluation (Bhatt, Weller, and Moura 2020).
| Model | Comprehensiveness | Sufficiency |
| 0.75 | 0.50 | |
| 0.90 | 0.35 |
@@
We expect these findings to extend to deep neural networks as well. In these more complex models, various components, such as activation functions and attention mechanisms, induce feature interactions (Tsang, Rambhatla, and Liu 2020).
| Model | Comprehensiveness | Sufficiency |
| 0.6 | 0.325 | |
| 0.925 | 0.65 |
@@
uses exhaustive search to find the exact lower and upper AOPC limits. It calculates the AOPC score for all possible feature orderings , where is the number of features. While precise, NAOPC’s time complexity makes it prohibitively slow for high-dimensional inputs.
@@
Experimental Setup
Our experiments address three key questions: 1) Do AOPC lower and upper limits vary across deep neural network models? 2) How does NAOPC affect model faithfulness rankings? and 3) How accurately does NAOPC approximate NAOPC? This section outlines the experimental designs, including the models, datasets, and feature attribution methods employed to investigate these questions.
@@
Our study focuses on six pre-trained sentiment classification models, all publicly available on Huggingface (Morris et al. 2020). These models are fine-tuned versions of BERT and RoBERTa: BERT, BERT, BERT, RoBERTa, RoBERTa, and RoBERTa.
These models were specifically chosen for their use in previous studies comparing AOPC scores across models (Hase, Xie, and Bansal 2021; Li et al. 2023; Bhalla, Srinivas, and Lakkaraju 2023; Chrysostomou and Aletras 2022). The models differ in two key aspects:
@@
Training Dataset: Each model was trained on either Yelp, IMDB, or SST2 (Zhang, Zhao, and LeCun 2015; Maas et al. 2011; Socher et al. 2013), enabling investigation of how dataset-specific characteristics affect AOPC score limits. The suffix in each model’s name (e.g., BERT) indicates the dataset on which it was trained.
| # examples | # words per example | |
| Yelp | 339 | 5 (4–7) |
| SST2 | 66 | 8 (7–9) |
| Yelp | 1000 | 52 (30–92) |
| SST2 | 400 | 19 (13–26) |
| IMDB | 1000 | 132 (98–173) |
Data
We evaluate these models using subsets of the same three datasets: Yelp, IMDB, and SST2. These datasets were chosen for their prevalence in cross-model AOPC score comparison studies (Hase, Xie, and Bansal 2021; Bhalla, Srinivas, and Lakkaraju 2023; Li et al. 2023).
To address varying computational requirements and ensure comprehensive analysis, we create short-sequence and long-sequence subsets from each dataset’s test set111For SST2, we use the validation set as our test set, as its test set is unlabeled.. Table 3 summarizes the key statistics of our subsets. The short subsets, Yelp and SST2, contain examples with up to 12 features, enabling computationally intensive evaluations such as NAOPC. We do not include an IMDB subset due to the scarcity of examples with fewer than 12 features in this dataset.
For the long subsets, Yelp and IMDB, we randomly sample 1000 examples from each dataset. We choose this sample size to balance computational feasibility with the need for a statistically significant sample size. We exclude examples exceeding 512 tokens due to model constraints. SST2 includes all 400 available test examples, as this was already a manageable size.
@@
We implement seven feature attribution methods: two transformer-specific, three gradient-based, and three perturbation-based (see Lyu, Apidianaki, and Callison-Burch (2024) for an extensive overview of feature attribution methods). The transformer-specific methods, Attention (Jain and Wallace 2019) and DecompX (Modarressi et al. 2023) are specifically designed for transformer architectures. Attention calculates the feature attribution scores only using the attention weights in the final layer, while DecompX uses all the components and layers in the transformer architecture. The gradient-based methods, InputXGrad (Sundararajan, Taly, and Yan 2017), Integrated Gradients (Sundararajan, Taly, and Yan 2017), and Deeplift (Shrikumar, Greenside, and Kundaje 2017), use backpropagation to quantify the influence of input features on output. The perturbation-based methods, LIME (Ribeiro, Singh, and Guestrin 2016), KernelSHAP (Lundberg and Lee 2017), and Occlusion@1 (Ribeiro, Singh, and Guestrin 2016), assess the impact on output confidence by occluding input features.
In this paper, we use a perturbation function that replaces tokens with the mask token to calculate IG, Deeplift, LIME, KernelSHAP, and Occlusion@1. We also use this perturbation function to calculate the AOPC scores as recommended by Hase, Xie, and Bansal (2021).
@@
We aim to show that the lower and upper limits of the AOPC scores vary between the models and inputs. We do so by calculating each model’s upper and lower AOPC limits using an exhaustive search for each example in Yelp (same strategy used by NAOPC). We then compare the models’ lower and upper limit distributions to demonstrate their differences.
@@
To answer these questions, we compare the sufficiency and comprehensiveness scores using AOPC and NAOPC. Specifically, we compare AOPC with NAOPC for all possible pairs of models and feature attribution methods on the Yelp, IMDB, and SST2 datasets. For the Yelp and SST2 datasets, we compare AOPC with both NAOPC and NAOPC. We analyze the results to find whether NAOPC changes the ranking of which models and feature attribution methods are the most faithful. We use a beam size of 5 when calculating NAOPC.
@@
We aim to demonstrate that NAOPC is a fast and reliable approximation of NAOPC. First, we demonstrate that the faithfulness rankings produced with NAOPC and NAOPC are similar on Yelp and SST2. We cannot make this comparison on the three other dataset subsets because calculating NAOPC on high-dimensional inputs is prohibitively slow. Instead, we calculate NAOPC with increasing beam sizes and analyze the change in results.
Results
The lower and upper limits vary between models
@@ Results
Figure 1 shows significant variations in the distributions of lower and upper AOPC score limits across different models on the Yelp test set. Each model has a distribution rather than a single value because individual inputs also influence the AOPC limits for each model. The clear differences in these distributions across models highlight that direct comparisons of comprehensiveness and sufficiency scores between models can be misleading without proper normalization. Moreover, these variations make interpreting AOPC scores in absolute terms challenging. For instance, an AOPC score of 0.25 might be considered high for RoBERTa but low for BERT. These distribution shifts across models emphasize the need for NAOPC. In Appendix B, we provide distributions of lower and upper AOPC score limits across different models on the SST2.
NAOPC alters faithfulness rankings
@@
Normalization through NAOPC substantially altered models’ faithfulness rankings while preserving the relative performance of feature attribution methods. This effect is clearly visible in Figure 2, where lines of different colors (representing different models) frequently intersect between AOPC and NAOPC rankings across Yelp, IMDB, and SST2 datasets. In contrast, lines of the same color (representing feature attribution methods within a model) rarely cross, indicating stability in their relative rankings. The impact of normalization appears even more pronounced in shorter text datasets, as shown in Figure 3 for Yelp.
@@
These visual observations are quantitatively supported by the Kendall rank correlation coefficients presented in Table 4. For model comparisons, correlations between AOPC and NAOPC scores are notably lower for sufficiency (ranging from 0.43 to 0.72) than for comprehensiveness (0.87 to 0.93), confirming that normalization significantly impacts model rankings, especially for sufficiency metrics. In contrast, feature attribution methods show consistently high correlations (0.90 to 0.99) for both comprehensiveness and sufficiency across all datasets, underscoring their ranking stability under normalization. This pattern is consistent across all datasets.
| Dataset | Group | Comprehensiveness | Sufficiency |
| Yelp | Model | 0.87 | 0.47 |
| FA | 0.97 | 0.97 | |
| SST-2 | Model | 0.89 | 0.43 |
| FA | 0.90 | 0.92 | |
| IMDB | Model | 0.93 | 0.72 |
| FA | 0.99 | 0.97 |
NAOPC accurately approximates NAOPC
Our analysis demonstrates that NAOPC accurately approximates NAOPC across various dataset dimensions. For low-dimensional input examples, Figure 3 shows nearly identical rankings produced by NAOPC and NAOPC on Yelp. We provide similar results observed for SST2 in Appendix C. For high-dimensional inputs, Figure 4 reveals that increasing the beam size beyond 5 yields minimal changes in the identified lower and upper limits for RoBERTa and BERT on Yelp, while we show comparable outcomes for a broad range of models and datasets in Appendix D. Importantly, NAOPC with a beam size of 5 proves efficient for practical use, requiring an average of 2 minutes per example in Yelp on a single A100 GPU while maintaining accuracy.
Discussion
Should one always normalize the AOPC scores?

### v2

@@ Abstract
Deep neural network predictions are notoriously difficult to interpret. Feature attribution methods aim to explain these predictions by identifying the contribution of each input feature. Faithfulness, often evaluated using the area over the perturbation curve (AOPC), reflects feature attributions’ accuracy in describing the internal mechanisms of deep neural networks. However, many studies rely on AOPC to compare faithfulness across different models, which we show can lead to false conclusions about models’ faithfulness. Specifically, we find that AOPC is sensitive to variations in the model, resulting in unreliable cross-model comparisons. Moreover, AOPC scores are difficult to interpret in isolation without knowing the model-specific lower and upper limits. To address these issues, we propose a normalization approach, Normalized AOPC (NAOPC), enabling consistent cross-model evaluations and more meaningful interpretation of individual scores. Our experiments demonstrate that this normalization can radically change AOPC results, questioning the conclusions of earlier studies and offering a more robust framework for assessing feature attribution faithfulness. Our code is available at https://github.com/JoakimEdin/naopc.
Normalized AOPC: Fixing Misleading Faithfulness Metrics for Feature Attribution Explainability
Joakim Edin1,2∗ Andreas G. Motzfeldt 1,3∗ Casper L. Christensen1∗ Tuukka Ruotsalo2,4 Lars Maaløe1 Maria Maistro2 1Corti 2University of Copenhagen 3IT University of Copenhagen 4LUT University ∗Equal contributions {je,clu,amo}@corti.ai
1 Introduction
@@
Deep neural networks are often described as black boxes due to the difficulty in understanding their inner mechanisms Wei et al. (2022). This lack of interpretability can hinder their adoption in critical applications where trust is paramount Lipton (2018). For instance, if a diagnostic model predicts meningitis without an explanation, a physician confident in an influenza diagnosis might incorrectly dismiss the model’s prediction as an error.
@@
Feature attribution methods attempt to address this issue by quantifying each input feature’s contribution to a model’s output Danilevsky et al. (2020). In the meningitis example, such a method might identify “fever” and “stiff neck” as important features, potentially convincing the physician to reconsider the diagnosis. For these methods to be reliable, they must faithfully represent the model’s underlying reasoning process, avoiding misleading interpretations Jacovi and Goldberg (2020).
The Area Over the Perturbation Curve (AOPC) has become a standard metric for approximating faithfulness, with two main variants: sufficiency and comprehensiveness DeYoung et al. (2020); Lyu et al. (2024). However, we uncover a critical weakness: the minimum and maximum possible AOPC scores vary significantly across different models and inputs. For the same dataset, we found one model’s upper limit averaged 0.3 across examples while another averaged 0.8. These varying lower and upper limits of AOPC stem from how models transform inputs into outputs, specifically, how many input features each model uses and how it combines these features through interactions to produce predictions.
This weakness invalidates AOPC comparisons across models, affecting several studies in explainable AI. We found eleven studies in top machine learning venues that compare AOPC across models. These studies are spread among the following research directions: learning to explain Resck et al. (2024); Liu et al. (2022), developing self-explanatory model architectures Sekhon et al. (2023), making models more interpretable through training strategies (e.g., adversarial robustness) Bhalla et al. (2023); Li et al. (2023); Chen and Ji (2020); Chrysostomou and Aletras (2021); Xie et al. (2024), and analyzing explanation methods in different settings (e.g., out-of-distribution data) Chrysostomou and Aletras (2022); Hase et al. (2021); Nielsen et al. (2023). Our findings suggest these studies’ findings may be misguiding.
The varying limits also make AOPC scores difficult to interpret. For instance, is an AOPC score of 0.25 high? It is high if the upper limit is 0.3 but not if it is 0.8. Interpretable scores would help researchers identify models and inputs where explanation methods produce unfaithful explanations, enabling them to analyze these cases and debug their methods. Without knowing the model and input-specific limits, this systematic improvement of explanation methods remains difficult.
@@
We demonstrate that the minimum and maximum possible AOPC scores vary significantly across different models and inputs, which makes cross-model comparisons and isolated score interpretations problematic.
@@
We show empirically, with five datasets, four model architectures, and three NLP tasks, how NAOPC alters faithfulness rankings, highlighting the need to re-evaluate conclusions in previous studies about model faithfulness.
To facilitate adoption of these methods, we release AOPC, NAOPC, and NAOPC as a PyPI package111Our names are mentioned in the PyPI package. We will include a link in the camera-ready version..
2 Problem formulation
Area Over the Perturbation Curve (AOPC) measures the change in model output as input features are sequentially perturbed Samek et al. (2016). The perturbation can either remove, insert, or replace a feature with some pre-defined value. The final score is the average output change across all perturbation steps. Formally, the AOPC is calculated as follows:
where is a model, is an input vector with number of features, is the order to perturb the input features, and , is the perturbation function that removes, inserts or replaces the features in that are in . AOPC is used to calculate sufficiency and comprehensiveness as follows.
Comprehensiveness estimates the faithfulness of feature attribution scores by perturbing the input features in decreasing order, starting from the feature with the highest score DeYoung et al. (2020). A high comprehensiveness indicates that the features that are the highest ranked, according to the feature attributions, are important for the model’s output (i.e., higher is better). Comprehensiveness is calculated as follows:
@@ 2 Problem formulation
Sufficiency perturbs the input features in increasing order, starting from the feature with the lowest score DeYoung et al. (2020). In other words, sufficiency is comprehensiveness when flipping the feature ordering . A low sufficiency indicates that the lowest ranked features, according to the feature attributions, are irrelevant to the model output (i.e., lower is better). Sufficiency is calculated as follows:
@@
2.1 Models influence AOPC scores
Ideally, sufficiency and comprehensiveness should only measure the feature attribution’s faithfulness. However, in this section, we demonstrate that a model’s reasoning process heavily influences these AOPC metrics. Specifically, we show with two toy examples that 1) the more features a model relies on, the worse the sufficiency and comprehensiveness scores, and 2) a model’s features interactions impact the best possible comprehensiveness and sufficiency scores.
@@ 2.1 Models influence AOPC scores
The two models use the same architecture, but their parameter values differ. relies heavily on feature , while relies on more features. Different training strategies, data, or randomness can cause such model differences Hase et al. (2021).
Given an input vector , both models output . Since the models are linear, we can calculate the ground truth feature attributions by multiplying each input feature by its parameter. For , this results in for , and for . As these attributions represent the ground truths, one might expect the models to yield equal sufficiency and comprehensiveness scores. However, this is not the case. As shown in Table 1, achieves drastically worse sufficiency and comprehensiveness scores than simply because it relies on more features. While relying on fewer features could be a desirable model property, it should be measured using entropy instead of influencing the faithfulness evaluation Bhatt et al. (2020).
Model Comprehensiveness Sufficiency 0.75 0.50 0.90 0.35
@@
We expect these findings to extend to deep neural networks as well. In these more complex models, various components, such as activation functions and attention mechanisms, induce feature interactions Tsang et al. (2020).
Model Comprehensiveness Sufficiency 0.6 0.325 0.925 0.65
@@ 3 Normalized AOPC
uses an exhaustive search to find the exact lower and upper AOPC limits. It calculates the AOPC score for all possible feature orderings , where is the number of features. While precise, its time complexity makes it prohibitively slow for high-dimensional inputs.
@@
To select an appropriate beam size for NAOPC, we check if the upper and lower AOPC limits remain stable as we increase the beam size Freitag and Al-Onaizan (2017). We start with a beam size of 1 and double the beam size until the limits do not change more than a pre-selected threshold two iterations in a row.
4 Experimental Setup
Our experiments address three key questions: 1) Do AOPC lower and upper limits vary across deep neural network models? 2) How does NAOPC affect model faithfulness rankings? and 3) How accurately does NAOPC approximate NAOPC? This section outlines the experimental designs, including the datasets, models, and feature attribution methods employed to investigate these questions.
Data
Our experiments use five datasets: Yelp, IMDB, SST2, AG-News, and SNLI. Yelp, IMDB, and SST2 are sentiment classification datasets, which we chose for their prevalence in cross-model AOPC score comparison studies Hase et al. (2021); Bhalla et al. (2023); Li et al. (2023). AG-News is a text classification dataset, and SNLI is a natural language inference dataset Zhang et al. (2015); MacCartney and Manning (2008). We include these two datasets to evaluate whether our findings generalize to other tasks than sentiment classification.
To address varying computational requirements and ensure comprehensive analysis, we create short-sequence and long-sequence subsets from each dataset’s test set222For SST2, we use the validation set as our test set, as its test set is unlabeled.. Table 3 summarizes the key statistics of our subsets. The short subsets, Yelp and SST2, contain examples with up to 12 features, enabling computationally intensive evaluations such as NAOPC.
We create five long-sequence subsets: SST2, Yelp, IMDB, AG-News, and SNLI. We randomly sample 1000 examples from each dataset except SST2 (SST2 only comprises 400 examples). We choose this sample size to balance computational feasibility with the need for a statistically significant sample size. We exclude examples exceeding 512 tokens due to model constraints.
# examples # words per example Yelp 339 5 (4–7) SST2 66 8 (7–9) Yelp 1000 52 (30–92) SST2 400 19 (13–26) IMDB 1000 132 (98–173) AG-News 1000 37 (31–42) SNLI 1000 20 (16–26)
@@ 4 Experimental Setup
Our experiments use twelve language models, all publicly available on Huggingface Morris et al. (2020). These models differ in two key aspects:
@@
Training Dataset: Each model was trained on either Yelp, IMDB, SST2, AG-News, or SNLI Zhang et al. (2015); Maas et al. (2011); Socher et al. (2013), enabling investigation of how dataset-specific characteristics affect AOPC score limits. The lowercase suffix in each model’s name (e.g., BERT) indicates the dataset on which it was trained.
We provide an overview of the models in Table 5.
@@
We implement eight feature attribution methods: two transformer-specific, three gradient-based, and three perturbation-based (see Lyu et al. (2024) for an extensive overview of feature attribution methods). The transformer-specific methods, Attention Jain and Wallace (2019) and DecompX Modarressi et al. (2023) are specifically designed for transformer architectures. Attention calculates the feature attribution scores only using the attention weights in the final layer, while DecompX uses all the components and layers in the transformer architecture. The gradient-based methods, InputXGrad Sundararajan et al. (2017), Integrated Gradients Sundararajan et al. (2017), and Deeplift Shrikumar et al. (2017), use backpropagation to quantify the influence of input features on output. The perturbation-based methods, LIME Ribeiro et al. (2016), KernelSHAP Lundberg and Lee (2017), and Occlusion@1 Ribeiro et al. (2016), assess the impact on output confidence by occluding input features.
In this paper, we use a perturbation function that replaces tokens with the mask token to calculate IG, Deeplift, LIME, KernelSHAP, and Occlusion@1. We also use this perturbation function to calculate the AOPC scores as recommended by Hase et al. (2021). We used the end-of-sequence token for GPT-2 because it does not support mask tokens nor pad tokens.
@@
We aim to show that the lower and upper limits of the AOPC scores vary between the models and inputs. We do so by calculating each model’s upper and lower AOPC limits using an exhaustive search for each example in Yelp (same search strategy used by NAOPC). We then compare the models’ lower and upper limit distributions to demonstrate their differences.
@@
To answer these questions, we compare the sufficiency and comprehensiveness scores using AOPC and NAOPC. Specifically, we compare AOPC with NAOPC for all possible pairs of models and feature attribution methods on the long-sequence datasets. For the short-sequence datasets, we compare AOPC with both NAOPC and NAOPC. We analyze the results to find whether NAOPC changes the ranking of which models and feature attribution methods are the most faithful. We use a beam size of 5 when calculating NAOPC on all datasets except for AG-News, which required a beam size of 1000.
@@
We aim to demonstrate that NAOPC is a fast and reliable approximation of NAOPC. With a sufficiently large beam size, NAOPC is equivalent to NAOPC. The question is if NAOPC can accurately approximate NAOPC with small beam sizes.
@@
First, we demonstrate that the faithfulness rankings produced with NAOPC and NAOPC are similar on Yelp and SST2 with a beam size of 5. We cannot make this comparison on the five long-sequence dataset subsets because calculating NAOPC on high-dimensional inputs is prohibitively slow. Instead, we calculate NAOPC with increasing beam sizes and analyze the change of the lower and upper AOPC limits. If the AOPC limits stabilize at small beam sizes, it indicates that NAOPC can efficiently and reliably approximate NAOPC.
5 Results
5.1 The lower and upper limits vary between models
@@
Figure 1 shows significant variations in the distributions of lower and upper AOPC score limits across different models on the Yelp test set. Each model has a distribution rather than a single value because individual inputs also influence the AOPC limits for each model. The clear differences in these distributions across models highlight that direct comparisons of AOPC scores between models can be misleading without proper normalization. Moreover, these variations make interpreting AOPC scores in absolute terms challenging. Figure 1 depicts an upper limit of around 0.3 for RoBERTa and 0.8 for BERT, therefore an AOPC score of 0.25 might be considered high for RoBERTa but low for BERT. These distribution shifts across models emphasize the need for NAOPC. Figure 5 depicts a similar pattern on SST2.
5.2 NAOPC alters faithfulness rankings
@@
Normalization through NAOPC substantially altered models’ faithfulness rankings while preserving the relative performance of feature attribution methods. This effect is clearly visible in Figure 2, where lines of different colors (representing different models) frequently intersect between AOPC and NAOPC rankings across Yelp, IMDB, and SST2 datasets. In contrast, lines of the same color (representing feature attribution methods within a model) rarely cross, indicating stability in their relative rankings. We see a similar trend on AG-News and SNLI in Figure 7. The impact of normalization appears even more pronounced in shorter text datasets, as shown in Figure 3 for Yelp.
@@
These visual observations are quantitatively supported by the Kendall rank correlation coefficients presented in Table 4 Kendall (1948). Correlations between AOPC and NAOPC scores are notably lower for the model comparisons than for the feature attribution comparisons. This pattern is consistent across all datasets.
Dataset Group Comp Suff Yelp Model 0.87 0.47 FA 0.97 0.97 SST-2 Model 0.89 0.43 FA 0.90 0.92 IMDB Model 0.93 0.72 FA 0.99 0.97 SNLI Model 0.71 1.0 FA 0.81 0.86 AG-News Model 0.25 0.67 FA 0.90 0.83
5.3 NAOPC accurately approximates NAOPC
Our analysis demonstrates that NAOPC accurately approximates NAOPC across various dataset dimensions. For low-dimensional input examples, Figure 3 shows nearly identical rankings produced by NAOPC and NAOPC on Yelp. Figure 6 depicts similar results for SST2.
For RoBERTa and BERT on Yelp, Figure 4 shows that a beam size of 5 is sufficient for stable results. However, the same figure demonstrates that BERT requires a substantially larger beam size. Figures 8, 9 and 10 confirm this pattern across all datasets, with AG-News being the only dataset requiring a larger beam size. We explore the reasons for this behavior and its implications in the next section.
6 Discussion
6.1 Does NAOPC require too much compute to be practically useful?
NAOPC is computationally more intensive than AOPC. Computing AOPC requires forward passes, NAOPC requires forward passes, and NAOPC requires forward passes, where is the number of input features, and is the beam size. While the exponential complexity of NAOPC makes it impractical for most inputs, our experiments demonstrate that NAOPC is feasible in many scenarios.
The computational cost of NAOPC depends on beam size, input length, and model inference time. With a small beam size (B=5), which suffices for many datasets, computing NAOPC for BERT (110 million parameters) on a hundred-feature example took around one minute on an A100 GPU. While requirements grow quadratically with input length, we found this manageable for several hundred tokens: processing 512-token inputs (BERT’s maximum) took approximately 10 minutes per example, averaged over 100 examples. Since NAOPC scales linearly with model size, larger models remain feasible to evaluate. Importantly, normalization factors only need to be computed once per model-dataset pair and can be reused, significantly reducing overall cost.
However, some datasets require larger beam sizes for accurate normalization. For instance, models trained on AG-News required a beam size of 1,000 to achieve stable results. This requirement does not appear to relate to input length, as AG-News comprises shorter sequences than Yelp and SST2. We speculate this might be due to feature interactions where multiple features must be removed together to measure their true impact on the model’s prediction. In such cases, a larger beam size is necessary to ensure these feature combinations are explored during the search. However, further research is needed to verify this hypothesis and understand what drives beam size requirements.
To help researchers assess requirements upfront, we provide tools for estimating necessary beam sizes for specific model-dataset combinations, allowing evaluation of cross-model comparison feasibility given computational constraints.
6.2 Should one always normalize the AOPC scores?
Given that computing, NAOPC requires additional computational resources, a natural question arises: when is this extra computation necessary? Our findings demonstrate that normalization is essential in two scenarios: comparing AOPC scores across different models and interpreting individual scores in relation to a model’s theoretical limits.